OpenStudy (anonymous):

The drama club sells 200lb of fruit to raise money. The fruit is sold in 5lb bags and 10lb bags. What is an equation to represent the number of each type of bag that the club should sell. pleease help! :(

OpenStudy (agent0smith):

x: number of 5 lb bags y: number of 10 lb bags They sell 200 lb total... so 5x + 10y = 200

OpenStudy (anonymous):

and then to simplify into y=mx+b... would it be y = -1/2x + 20 ?

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

awesome thank you so much!